Nerio is now 2 years younger than Jean, while Angel is one year younger than Mon. If the sum of Mon's Age in 3 years, Jean's age 2 years ago, and Angel's age next year is 78, find their present ages.
Thanks!

Answer by ankor@dixie-net.com(22740)

Nerio is now 2 years younger than Jean,
n = j-2
:
while Angel is one year younger than Mon.
a = m-1
:
If the sum of Mon's Age in 3 years, Jean's age 2 years ago, and Angel's age next year is 78,
(m+3) + (j-2) + (a+1) = 78
m + j + a + 4 = 78
m + j + a = 78 - 4
m + j + a = 74
m + j + (m-1) = 74
2m + j = 74 + 1
2m + j = 75
:
This about all you can do with this, Where does Nerio appear in this problem?




find their present ages.